I need help understanding the steps to answer this problem. Thanks!

A voltaic cell is constructed that uses the following reaction and operates at 298 K:
Zn(s)+Ni2+(aq)→Zn2+(aq)+Ni(s).

1.) What is the emf of this cell under standard conditions?

2.) What is the emf of this cell when [Ni2+]= 2.80 M and [Zn2+]= 0.130 M ?

3.) What is the emf of the cell when [Ni2+]= 0.150 M and [Zn2+]= 0.999 M ?

Express your answer using two significant figures.
Answer #1

1.

Anode reaction

Zn -2e  \rightarrow Zn2+ , Eo(Zn2+/Zn ) = - 0.76 V

Cathode reaction

Ni2+ +2e \rightarrow Ni , Eo(Ni2+/Ni) = - 0.25 V

So emf at standard condition

Eocell = Eocathode - Eoanode = - 0.25 - (-0.76) = 0.51 V

The cell reaction is

Zn(s) + Ni2+(aq) \to Ni​​​​​​ (s) + Zn2+(aq)

Number of electron transferred (n) = 2

Now, nernst equation for the cell reaction at 298 K is

For, Solids, [Zn] =[Ni] = 1

Ecell = Eocell - 0.059 log [Zn2+]/[Ni2+] (1)

Or, Ecell = 0.51 - 0.0295 * log[Zn2+]/[Ni2+] (2)

2)

When, [Zn2+] = 0.130 M, [Ni2+]= 2.80 M

Or, Ecell = 0.51 - 0.059 log (0.130/2.80)

Or, Ecell = 0.51 - 0.0295*(-1.33)

Or, Ecell = 0.51 + 0.04 = 0.55 V.( In two significant figure)

3)

[Ni2+] = 0.150 M, [Zn2+]= 0.999 M

Then, using Eq.2

Ecell = 0.51 - 0.0295* log ( 0.999/0.150)

Ecell = 0 51 - 0.0295 * (0.823)

Or, Ecell = 0.51 - 0.024 = 0.486 V = 0.49 V ( in two significant figure)